OpenStudy (anonymous):

A movie rental program charges 10 per month plus an additional 2 for each movie rented. Which equation could be used to determine how many movies were rented if your monthly bill was 24 ?

OpenStudy (anonymous):

C(x)=2x+10 =24 solve for x for # of movies

OpenStudy (calculusxy):

24=(10+2)m

OpenStudy (calculusxy):

10+2 is the amount of money

OpenStudy (anonymous):

"24=(10+2)m" what part represents the number of movies?

OpenStudy (calculusxy):

m=movies

OpenStudy (anonymous):

1 = 12 2 = 24 3 = 36?

OpenStudy (anonymous):

24=(10+2)m

OpenStudy (calculusxy):

@KeeMoney Wasn't that what I wrote?

OpenStudy (anonymous):

that equation charges a monthly fee and a movie fee every time a movie is rented.....

OpenStudy (anonymous):

it doesn't flutter matter, it's wrong

OpenStudy (anonymous):

@calculusxy does it matter ?

OpenStudy (anonymous):

thanks guys

OpenStudy (calculusxy):

Yes, the 10 is the monthly fee and the 2 is the additional fee. The total will be 12 per moth.

OpenStudy (calculusxy):

@KeeMoney I don't want to argue or anything. But you can't put the same answer as to someone unless you are mentioning that person from whom you copied.

OpenStudy (anonymous):